# An Efficient Machine Learning Framework for Option Pricing via Fourier Transform

### v1

@@ Abstract
The increasing need for rapid recalibration of option pricing models in dynamic markets places stringent computational demands on data generation and valuation algorithms. In this work, we propose a hybrid algorithmic framework that integrates the smooth offset algorithm (SOA) with supervised machine learning models for the fast pricing of multiple path-independent options under exponential Lévy dynamics. Building upon the SOA-generated dataset, we train neural networks, random forests, and gradient boosted decision trees to construct surrogate pricing operators. Extensive numerical experiments demonstrate that, once trained, these surrogates achieve order-of-magnitude acceleration over direct SOA evaluation. Importantly, the proposed framework overcomes key numerical limitations inherent to fast Fourier transform–based methods, including the consistency of input data and the instability in deep out-of-the-money option pricing.
@@ 1 Introduction
Compared with traditional algorithms, machine learning (ML) has seen growing adoption in option pricing. Recent research increasingly leverages ML models to establish the mapping between option attributes and their theoretical prices. For example, [15] investigates ordinary least squares, radial basis function networks, multi-layer perceptron networks, and projection pursuit on the Black–Scholes option pricing problem. It demonstrates that ML models achieve accurate pricing and delta-hedging. Likewise, [11] analyzes basket call options. It shows that neural networks (NNs) can compute prices at a dramatically faster speed. This speed is approximately one million times faster than traditional methods, while maintaining satisfactory accuracy. To address the problems of inadequate model accuracy and anti-noise stability in option pricing, [17] adopts an experimental evaluation strategy integrated with financial theories, investigates the application effect of ensemble learning and the interaction between sliding windows and noise, verifies the applicability of ensemble learning to option pricing, and clarifies the adaptation logic between the two as well as the influence law on the pricing stability of the model. Other ML models address option pricing problems by numerically solving the associated PDEs through NNs. This method is commonly referred to as physics informed neural networks (PINNs) ([10, 12, 21, 26]). Although traditional machine learning models have achieved certain breakthroughs in pricing accuracy and efficiency, they still face core challenges in practical financial applications. The dynamic changes in market conditions require models to have rapid updating capabilities, while pricing models typically rely on large-scale data for training. Traditional numerical methods exhibit inefficiencies in data generation, and illiquid option contracts also suffer from data scarcity, making it difficult to support high-frequency iterations. Thus, we propose a pricing algorithm integrating SOA with machine learning. This algorithm uses option data generated by SOA as the training set, allowing the data-driven model to learn the mapping relationship between option attributes and their corresponding prices. Numerical experiments demonstrate that compared with fast Fourier transform (FFT) the machine learning models trained in this way is not constrained by input conditions and maintains good numerical stability.
@@ 3.1 Revised CMA
CMA modifies the original option pricing formula by introducing an offset term, ensuring that the FT of the resulting expression exists and admits a close-form representation. Consider an European or digital option whose underlying stock price $\left\{S_{t}\right\}_{0\leqslant t\leqslant T}$ follows exponential Lévy process with $S_{t}=\exp\left(rt+X_{t}^{\dagger}\right)$. The price function modified by the CM offset term is
@@
Consider an European or digital option whose underlying stocks price $\left\{S_{t}\right\}_{0\leqslant t\leqslant T}$ follows exponential Lévy process with $S_{t}=\exp\left(rt+X_{t}^{\dagger}\right)$. The price function modified by the smooth offset term is
@@ 3.2 Numerical Experiments for SOA and CMA
where the weights $w_{j}$ are the weight coefficient corresponding to the node in Simpson’s numerical integration rule.
@@ 4.1 FFT-based Framework
In this section, we integrate SOA with the fast Fourier transform (FFT). It is noteworthy that although FFT is also employed in [9], their formulation is designed for single option pricing tasks, whereas we generalizes the FFT-based algorithm to handle multiple options simultaneously. In detail, it is natural to consider the application of FFT for multiple-option pricing tasks, as these tasks involve evaluating a set of summations of the form in (3.4). However, the framework is applicable only when all options share identical attributes, with strike prices being the sole varying attribute. Under this condition, pricing the entire option portfolio becomes equivalent to computing a discrete Fourier transform (DFT), which can be efficiently accelerated by FFT.
@@ Acknowledgments
The authors would like to express their appreciation to the referees for their useful comments and the editors. Liying Zhang is supported by the National Natural Science Foundation of China (No. 11601514 and No. 11971458), the Fundamental Research Funds for the Central Universities (No. 2023ZKPYL02 and No. 2023JCCXLX01) and the Yueqi Youth Scholar Research Funds for the China University of Mining and Technology-Beijing (No. 2020YQLX03).

### v2

@@ Abstract
The increasing need for rapid recalibration of option pricing models in dynamic markets places stringent computational demands on data generation and valuation algorithms. In this work, we propose a hybrid algorithmic framework that integrates the smooth offset algorithm (SOA) with supervised machine learning models for the fast pricing of multiple path-independent options under exponential Lévy dynamics. Building upon the SOA-generated dataset, we train neural networks, random forests, and gradient boosted decision trees to construct surrogate pricing operators. Extensive numerical experiments demonstrate that, once trained, these surrogates achieve order-of-magnitude acceleration over direct SOA evaluation. Importantly, the proposed framework overcomes key numerical limitations inherent to fast Fourier transform–based methods, including input data consistency and instability in deep out-of-the-money option pricing.
@@ 1 Introduction
Compared with traditional algorithms, machine learning (ML) has seen growing adoption in option pricing. Recent research increasingly leverages ML models to establish the mapping between option attributes and their theoretical prices. For example, [15] investigates ordinary least squares, radial basis function networks, multi-layer perceptron networks, and projection pursuit on the Black–Scholes option pricing problem. It demonstrates that ML models achieve accurate pricing and delta-hedging. Likewise, [11] analyzes basket call options. It shows that neural networks (NNs) can compute prices at a dramatically faster speed. This speed is approximately one million times faster than traditional methods, while maintaining satisfactory accuracy. To address the problems of inadequate model accuracy and anti-noise stability in option pricing, [17] adopts an experimental evaluation strategy integrated with financial theories, investigates the application effect of ensemble learning and the interaction between sliding windows and noise, verifies the applicability of ensemble learning to option pricing, and clarifies the adaptation logic between the two as well as the influence law on the pricing stability of the model. Other ML models address option pricing problems by numerically solving the associated PDEs through NNs. This method is commonly referred to as physics informed neural networks (PINNs) ([10, 12, 21, 26]). Although traditional machine learning models have achieved certain breakthroughs in pricing accuracy and efficiency, they still face core challenges in practical financial applications. The dynamic changes in market conditions require models to have rapid updating capabilities, while pricing models typically rely on large-scale data for training. Traditional numerical methods exhibit inefficiencies in data generation, and illiquid option contracts also suffer from data scarcity, making it difficult to support high-frequency iterations. Thus, we propose a pricing algorithm integrating SOA with machine learning. This algorithm uses option data generated by SOA as the training set, allowing the data-driven model to learn the mapping relationship between option attributes and their corresponding prices. Numerical experiments demonstrate that compared with fast Fourier transform (FFT) the machine learning models trained in this way are not constrained by input conditions and maintains good numerical stability.
@@ 3.1 Revised CMA
CMA modifies the original option pricing formula by introducing an offset term, ensuring that the FT of the resulting expression exists and admits a closed-form representation. Consider a European or digital option whose underlying stock price $\left\{S_{t}\right\}_{0\leqslant t\leqslant T}$ follows exponential Lévy process with $S_{t}=\exp\left(rt+X_{t}^{\dagger}\right)$. The price function modified by the CM offset term is
@@
Consider a European or digital option whose underlying stock price $\left\{S_{t}\right\}_{0\leqslant t\leqslant T}$ follows exponential Lévy process with $S_{t}=\exp\left(rt+X_{t}^{\dagger}\right)$. The price function modified by the smooth offset term is
@@ 3.2 Numerical Experiments for SOA and CMA
where the weights $w_{j}$ are the weight coefficients corresponding to the node in Simpson’s numerical integration rule.
@@ 4.1 FFT-based Framework
In this section, we integrate SOA with the fast Fourier transform. It is noteworthy that although FFT is also employed in [9], their formulation is designed for single option pricing tasks, whereas we generalize the FFT-based algorithm to handle multiple options simultaneously. In detail, it is natural to consider the application of FFT for multiple-option pricing tasks, as these tasks involve evaluating a set of summations of the form in (3.4). However, the framework is applicable only when all options share identical attributes, with strike prices being the sole varying attribute. Under this condition, pricing the entire option portfolio becomes equivalent to computing a discrete Fourier transform (DFT), which can be efficiently accelerated by FFT.
@@ Acknowledgments
The authors would like to express their appreciation to the referees for their useful comments and the editors. Liying Zhang is supported by the National Natural Science Foundation of China (No. 11601514 and No. 11971458), the Fundamental Research Funds for the Central Universities (No. 2023ZKPYL02 and No. 2023JCCXLX01) and the Yueqi Youth Scholar Research Funds for the China University of Mining and Technology-Beijing (No. 2020YQLX03). 2025 Basic Sciences Initiative in Mathematics and Physics.